# thanks: Equal contributionthanks: Equal contribution

### 23 Mar 2026 v1

@@ Abstract
While our earlier short communication [45] described the fabrication protocol and basic swimming behaviour of chlamylipo, a comprehensive physical picture is still lacking. To understand the underlying physics and to establish a simple deformation-velocity equation that is valid over a broad parameter space, we varied the experimental conditions to observe a wide variety of swimming modes. The movements and deformations of chlamylipo are observed under six distinct conditions by modifying the osmotic condition of the outer solution (hypotonic, isotonic, and hypertonic) and using two different Chlamydomonas strains (wildtype CC-125 and oda1). The osmotic condition was varied to alter the reduced volume of the liposome, as discussed in detail later. The flagella-mutant strain oda1 is known for its slower beating of flagella, and its beating frequency and swimming velocity are nearly half of those of WT (CC-125) [30, 14]. A schematic in Fig. 2a represents the obtained parameters via image processing; the swimming velocity $U=2.10\pm 2.55$ $\mathrm{\mu m}/s$, the beating frequency $f=23.6\pm 13.8$ Hz, the liposome radius $R_{\mathrm{L}}=7.73\pm 1.75$ $\mathrm{\mu m}$, length of the major axis of Chlamydomonas $R_{\mathrm{C}}=3.72\pm 0.73$ $\mathrm{\mu m}$, and the protrusion $p=0.90\pm 0.45$ $\mathrm{\mu m}$ are obtained from the entire dataset $N=487$ with all six conditions. The protrusion $p=P-R_{\mathrm{L}}$ is evaluated by subtracting the liposome radius $R_{\mathrm{L}}$ from the farthest distance $P$ from the centroid. The dimensionless protrusion is defined as
@@
and the value was $p^{*}=0.13\pm 0.07$; the result suggests that the protrusion $p$ is approximately $13\%$ of the liposome radius on average. Now, can parameters describing membrane deformation be used to explain the observed swimming velocity? Figure 2b shows the correlation matrix of the obtained variables. The same figure with the scatter plots is shown in Fig. LABEL:fig:S1. While the velocity $U$ does not correlate with the farthest protrusion $P$, it strongly correlates with the dimensionless protrusion $p^{*}$ (correlation $r=0.66$), weakly with $f$ and $R_{\mathrm{C}}$, and has weakly negative correlation with $R_{\mathrm{L}}$. Notably, as we will discuss later, there is a negative correlation between $p^{*}$ and $R_{\mathrm{L}}$ and a weak negative correlation between $p^{*}$ and $f$.
@@
Figures 4b–c and Supplemental Movie S4 show the chlamylipo movement under alternating UV and blue light pulses at every 20 s intervals; the irradiating time of each pulse was approximately 500 ms. Note that we used the green light to guide the chlamylipo towards the right side, UV and blue light to switch motility, and red light is used for observation, as Chlamydomonas exhibits no phototactic response to red light. The top panel of Fig. 4b shows the distance from the initial position $|\mathbf{x}(t)-\mathbf{x}(0)|$ where $\mathbf{x}(t)$ is the position at time $t$, while the bottom panel shows the swimming velocity $U$. Although the chlamylipo is in the non-motile state at the initial condition (i), it begins to swim forward quickly after the first UV pulse (ii) with a velocity $\approx 2.0$ $\mathrm{\mu m/s}$, and returns to the non-motile state right after the first blue-light pulse (iii). These motile and non-motile states can be switched reversibly and cyclically, and we have successfully repeated this clutch switching more than seven times, as shown in the figures. These two states enable precise manipulation of the swimmer’s motion; Fig. 4d and Supplemental Movie S5 show that we can trace letters “CL”, the initials of chlamylipo, by steering the position. The top image of Fig 4d shows the trajectory colored with the elapsed time, whereas the bottom image is colored with the swimming velocity, with marker radius inversely proportional to the local velocity magnitude. During tracing, after the swimmer reached each corner of CL letters, the chlamylipo was switched to the non-motile state by blue light; these pauses are indicated by large blue markers. The green light position was then shifted to set the next heading before switching back to the motile state. While steering is possible without the clutch, access to motile and non-motile states allowed precise cornering; pausing at each vertex of the “CL” trajectory enabled accurate reorientation and precise letter formation. Although this tracing is primarily a demonstration, the ability to pause the swimmer in place offers several practical applications and may prove valuable, as discussed in the conclusion. We also note that motility can be switched collectively by broadcasting the switching signal to all chlamylipos, or individually by selectively irradiating only the target chlamylipo, as shown in Fig. LABEL:fig:S2 and Supplemental Movie S6.
@@
Finally, we demonstrate that the chlamylipo can transport cargo within the membrane, and also release the inner cargo at an arbitrary position. As illustrated in Fig. 5a, the inner space of the liposome can serve as a compartment for encapsulating desired cargo [1]. The chlamylipo that co-encapsulates cargo can be fabricated by preparing an aqueous solution containing both Chlamydomonas and the target cargo during water-in-oil emulsion formation, and we successfully co-encapsulated micro-sized beads, or even Escherichia coli (data not shown). As shown in Fig. 5b and Supplemental Movie S7, the chlamylipo can propel even when co-encapsulated with 0.5 $\mathrm{\mu m}$ polystyrene beads. The swimming velocity of this chlamylipo is 2.89 $\mathrm{\mu m/s}$, which is comparable to that without cargo. After guiding the chlamylipo to the destination, the inner cargo can be released from the membrane using an NIR (near infrared) laser, as shown in Fig. 5c and Supplemental Movie S8. The snapshot images show a chlamylipo containing 1.0 $\mathrm{\mu m}$ magnetic beads (i) that was irradiated with an NIR laser (ii), leading to the release of the beads, which subsequently diffused away (iii). The magnetic beads were expelled because NIR-induced local heating increased the membrane fluidity, leading to transient pore formation and eventual membrane bursting [34, 49, 22]. Although our approach is constrained by the requirement that the cargos must be harmless to Chlamydomonas, such as micro-sized colloids or benign liquid cargo, we demonstrated that cargos can be co-encapsulated, transported to a target destination, and released on demand by NIR irradiation.
@@
In this study, we analysed the swimming of chlamylipo from a fluid-mechanical perspective, identified the factors that govern the swimming velocity, and proposed methods to adjust the velocity based on the understanding. A nondimensional analysis indicated that the speed scales with the beating frequency, the square of the protrusion amplitude, and the inverse of the liposome radius. The ratio of the Chlamydomonas radius to the liposome radius also plays an important role: a larger ratio allows the membrane to protrude farther, leading to faster swimming. Thus, if cargo capacity is not a priority and the goal is simply to obtain a faster chlamylipo, the Chlamydomonas-to-liposome size ratio should be set to a larger value. During the analysis, we also found that the mutant oda1, whose beating frequency and free-swimming velocity are about half those of the wild type, reaches a speed comparable to the wild type when encapsulated; the larger protrusion of oda1 appears to compensate for the lower frequency. The swimming velocity can be modulated by tuning the protrusion level in response to external stimuli, such as osmotic conditions and light switching. When the inner volume of the liposome increases due to the osmotic difference across the membrane, the chlamylipo swims poorly as the membrane has only limited degrees of freedom for deformation. When the inner volume decreases, on the other hand, the chlamylipo can swim faster thanks to the rich deformability of the membrane. By introducing the photoswitchable phospholipid AzoPC into the liposomal membrane, the motility can be reversibly controlled with light, and the membrane works as a “clutch”. The chlamylipo switches to a motile state after UV illumination as the effective membrane area increases, and it returns to a non-motile state after subsequent blue light exposure. Although the light-based control used here was a simple two-state switch, finer continuous tuning should be achievable in future work by more finely controlling the membrane area.

### 31 Mar 2026 v2

@@ Abstract
While our earlier short communication [45] described the fabrication protocol and basic swimming behaviour of chlamylipo, a comprehensive physical picture is still lacking. To understand the underlying physics and to establish a simple deformation-velocity equation that is valid over a broad parameter space, we varied the experimental conditions to observe a wide variety of swimming modes. The movements and deformations of chlamylipo are observed under six distinct conditions by modifying the osmotic condition of the outer solution (hypotonic, isotonic, and hypertonic) and using two different Chlamydomonas strains (wild-type CC-125 and mutant oda1). The osmotic condition was varied to alter the reduced volume of the liposome, as discussed in detail later. The flagella-mutant strain oda1 is known for its slower beating of flagella, and its beating frequency and swimming velocity are nearly half of those of WT (CC-125) [30, 14]. A schematic in Fig. 2a represents the obtained parameters via image processing; the swimming velocity $U=2.10\pm 2.55$ $\mathrm{\mu m}/s$, the beating frequency $f=23.6\pm 13.8$ Hz, the liposome radius $R_{\mathrm{L}}=7.73\pm 1.75$ $\mathrm{\mu m}$, length of the major axis of Chlamydomonas $R_{\mathrm{C}}=3.72\pm 0.73$ $\mathrm{\mu m}$, and the protrusion $p=0.90\pm 0.45$ $\mathrm{\mu m}$ are obtained from the entire dataset $N=487$ with all six conditions. The protrusion $p=P-R_{\mathrm{L}}$ is evaluated by subtracting the liposome radius $R_{\mathrm{L}}$ from the farthest distance $P$ from the centroid. The dimensionless protrusion is defined as
@@
and the value was $p^{*}=0.13\pm 0.07$; the result suggests that the protrusion $p$ is approximately $13\%$ of the liposome radius on average. Now, can parameters describing membrane deformation be used to explain the observed swimming velocity? Figure 2b shows the correlation matrix of the obtained variables. The same figure with the scatter plots is shown in Fig. S1. While the velocity $U$ does not correlate with the farthest protrusion $P$, it strongly correlates with the dimensionless protrusion $p^{*}$ (correlation $r=0.66$), weakly with $f$ and $R_{\mathrm{C}}$, and has weakly negative correlation with $R_{\mathrm{L}}$. Notably, as we will discuss later, there is a negative correlation between $p^{*}$ and $R_{\mathrm{L}}$ and a weak negative correlation between $p^{*}$ and $f$.
@@
Figures 4b–c and Supplemental Movie S4 show the chlamylipo movement under alternating UV and blue light pulses at every 20 s intervals; the irradiating time of each pulse was approximately 500 ms. Note that we used the green light to guide the chlamylipo towards the right side, UV and blue light to switch motility, and red light is used for observation, as Chlamydomonas exhibits no phototactic response to red light. The top panel of Fig. 4b shows the distance from the initial position $|\mathbf{x}(t)-\mathbf{x}(0)|$ where $\mathbf{x}(t)$ is the position at time $t$, while the bottom panel shows the swimming velocity $U$. Although the chlamylipo is in the non-motile state at the initial condition (i), it begins to swim forward quickly after the first UV pulse (ii) with a velocity $\approx 2.0$ $\mathrm{\mu m/s}$, and returns to the non-motile state right after the first blue-light pulse (iii). These motile and non-motile states can be switched reversibly and cyclically, and we have successfully repeated this clutch switching more than seven times, as shown in the figures. These two states enable precise manipulation of the swimmer’s motion; Fig. 4d and Supplemental Movie S5 show that we can trace letters “CL”, the initials of chlamylipo, by steering the position. The top image of Fig 4d shows the trajectory colored with the elapsed time, whereas the bottom image is colored with the swimming velocity, with marker radius inversely proportional to the local velocity magnitude. During tracing, after the swimmer reached each corner of CL letters, the chlamylipo was switched to the non-motile state by blue light; these pauses are indicated by large blue markers. The green light position was then shifted to set the next heading before switching back to the motile state. While steering is possible without the clutch, access to motile and non-motile states allowed precise cornering; pausing at each vertex of the “CL” trajectory enabled accurate reorientation and precise letter formation. Although this tracing is primarily a demonstration, the ability to pause the swimmer in place offers several practical applications and may prove valuable, as discussed in the conclusion. We also note that motility can be switched collectively by broadcasting the switching signal to all chlamylipos, or individually by selectively irradiating only the target chlamylipo, as shown in Fig. S2 and Supplemental Movie S6.
@@
Finally, we demonstrate that the chlamylipo can transport cargo within the membrane, and also release the inner cargo at an arbitrary position. As illustrated in Fig. 5a, the inner space of the liposome can serve as a compartment for encapsulating desired cargo [1]. The chlamylipo that co-encapsulates cargo can be fabricated by preparing an aqueous solution containing both Chlamydomonas and the target cargo during water-in-oil emulsion formation, and we successfully co-encapsulated micro-sized beads, or even Escherichia coli (data not shown). As shown in Fig. 5b and Supplemental Movie S7, the chlamylipo can propel even when co-encapsulated with 0.5 $\mathrm{\mu m}$ polystyrene beads. The swimming velocity of this chlamylipo is 2.89 $\mathrm{\mu m/s}$, which is comparable to that without cargo. After guiding the chlamylipo to the destination, the inner cargo can be released from the membrane using an NIR (near infrared) laser [46], as shown in Fig. 5c and Supplemental Movie S8. The snapshot images show a chlamylipo containing 1.0 $\mathrm{\mu m}$ magnetic beads (i) that was irradiated with an NIR laser (ii), leading to the release of the beads, which subsequently diffused away (iii). The magnetic beads were expelled because NIR-induced local heating increased the membrane fluidity, leading to transient pore formation and eventual membrane bursting [34, 50, 22]. Although our approach is constrained by the requirement that the cargos must be harmless to Chlamydomonas, such as micro-sized colloids or benign liquid cargo, we demonstrated that cargos can be co-encapsulated, transported to a target destination, and released on demand by NIR irradiation.
@@
In this study, we analysed the swimming of chlamylipo from a fluid-mechanical perspective, identified the factors that govern the swimming velocity, and proposed methods to adjust the velocity based on the understanding. A nondimensional analysis indicated that the speed scales with the beating frequency, the square of the protrusion amplitude, and the inverse of the liposome radius. The ratio of the Chlamydomonas radius to the liposome radius also plays an important role: a larger ratio allows the membrane to protrude farther, leading to faster swimming. Thus, if cargo capacity is not a priority and the goal is simply to obtain a faster chlamylipo, the Chlamydomonas-to-liposome size ratio should be set to a larger value. During the analysis, we also found that the mutant oda1, whose beating frequency and free-swimming velocity are about half those of the wild-type, reaches a speed comparable to the wild-type when encapsulated; the larger protrusion of oda1 appears to compensate for the lower frequency. The swimming velocity can be modulated by tuning the protrusion level in response to external stimuli, such as osmotic conditions and light switching. When the inner volume of the liposome increases due to the osmotic difference across the membrane, the chlamylipo swims poorly as the membrane has only limited degrees of freedom for deformation. When the inner volume decreases, on the other hand, the chlamylipo can swim faster thanks to the rich deformability of the membrane. By introducing the photoswitchable phospholipid AzoPC into the liposomal membrane, the motility can be reversibly controlled with light, and the membrane works as a “clutch”. The chlamylipo switches to a motile state after UV illumination as the effective membrane area increases, and it returns to a non-motile state after subsequent blue light exposure. Although the light-based control used here was a simple two-state switch, finer continuous tuning should be achievable in future work by more finely controlling the membrane area.